1. You count 12 cycles across 8.5 time divisions on an oscilloscope screen.
The time sweep is set to 45 micro seconds.
The voltage peak is measured to be 22 volts.
What is the frequency in HZ?
(Use three significant figures or N/A if no answer.)

2. You plot some data of Real Frequency versus HP Dial Frequency and get a linear trend of
y = 1.31 x + 230.
To what must you set the HP Dial Frequency in order to get a Real Frequency of 4270 [Hz]?
[Hz] (Use three significant figures or N/A if no answer.)

3. In a
Answer #1

1) freequency = 12 / (8.5 * 45* 10^-6) = 31.372 KHz

2) HP Dial Frequency = (real frequency -230 )/1.31 = 3083.970 HZ

3) We used a log-log plot because the relationship between the two variables is of  y=ax^k form which appears a straight line in Log-Log plot.

4) In mathematics, a Lissajous figure is the graph of a system of parametric equations

x=A\sin(at+\delta),\quad y=B\sin(bt),

The appearance of the figure is highly sensitive to the ratio a/b. For a ratio of 1, the figure is an ellipse, with special cases including circles (A = B, ? = ?/2 radians) and lines (? = 0). Another simple Lissajous figure is the parabola (a/b = 2, ? = ?/4).

Lissajous curves can also be generated using an oscilloscope (as illustrated). An octopus circuit can be used to demonstrate the waveform images on an oscilloscope. Two phase-shifted sinusoid inputs are applied to the oscilloscope in X-Y mode and the phase relationship between the signals is presented as a Lissajous figure.

On an oscilloscope, we suppose x is CH1 and y is CH2, A is amplitude of CH1 and B is amplitude of CH2, a is frequency of CH1 and b is frequency of CH2, so a/b is a ratio of frequency of two channels, finally, ? is the phase shift of CH1.

5) y frequency / x frequency = 2/3

--> 430/ x frequency = 2/3

--> x frequency = 645 Hz